If the demand and supply curve for T-Shirts at Nagindas Store are:





D = 100 - 6P, S = 28 + 3P





Where P is the price of T-Shirts in dollars ($), what is the quantity of T-Shirt bought and sold at





equilibrium?





Q

1
Expert's answer
2022-04-11T12:09:59-0400

"D=100-6P,"

"S=28+3P,"

"D=S,"

"100-6P=28+3P,"

"72=9P,"

"P=8,"

"D=S=52."